OpenStudy (anonymous):

Someone please help me! Write the equation 5x + y = 4 in slope-intercept form. A) y = -5x + 4 B) y = -5x - 4 C) y = 5x - 4 D) y = 5x + 4

OpenStudy (precal):

get x by itself how do you move 5x to the other side

OpenStudy (anonymous):

I'm not sure that's why I need help.

OpenStudy (precal):

when you move things to the opposite side take its opposite sign what sign is 5x

OpenStudy (anonymous):

I'm not sure if it would be positive or negative. I think the answer would be y = 5x + 4 but I'm not sure. I don't know if the 5x should be positive or negative.

OpenStudy (precal):

when a sign is not given or nothing is in front assume that it is positive what is the opposite of positive

OpenStudy (anonymous):

The opposite of positive is obviously negative..

OpenStudy (precal):

5x+y=4 -5x -5x y=4-5x

OpenStudy (precal):

or y =-5x+4

OpenStudy (anonymous):

So -5x + 4 would be the correct answer?

OpenStudy (precal):

yes but most important do you understand why

OpenStudy (anonymous):

Yes because when a sign is not given or nothing is in front of the number just assume that it is positive. Then move onto solving the problem by subtracting by the opposite of the number & that will give you your answer..

OpenStudy (precal):

good job :)

OpenStudy (anonymous):

Thank you for helping me.

OpenStudy (precal):

anytime